The product of 2 integers is -144. The sum of the integers is -7. What are the 2 integers?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

The product of 2 integers is -144. The sum of the integers is -7. What are the 2 integers?
-------------------
Try pairs of factors of 144.
Find 2 that differ by 7.

Answer by stanbon(75887)

The product of 2 integers is -144. The sum of the integers is -7.
What are the 2 integers?
--------
Equations:
xy = -144
x + y = -7
-----
Substitute for "x" and solve for "y"::
x = -7-y
-----
-(y+7)y = -144
y^2 + 7y - 144 = 0
------
Factor::
(y-9)(y+16) = 0
Positive solution:
y = 9
----
Solve for "x":
x + y = -7
x + 9 = -7
x = -16
-------------
Cheers,
Stan H.
-----------